OpenStudy (anonymous):

2.Solve. 5 = –(–z + 3) A. –8 B. 8 C. –2 D. 2

OpenStudy (anonymous):

im guessing this one is -2..

OpenStudy (goose.):

This one is a bit more tricky than your last one, with the (-z + 3) in parentheses... you must first distribute your - to each one inside, as it is basically -1 x -z and -1 x 3

OpenStudy (anonymous):

or plug in the numbers...

OpenStudy (goose.):

So that'll get you z - 3

OpenStudy (goose.):

Then just like last time, isolate your variable, add 3 to each side.

OpenStudy (anonymous):

5=(-2+3)?

OpenStudy (goose.):

5 + 3 = z - 3 + 3

OpenStudy (anonymous):

oh yeah

OpenStudy (goose.):

You know -3 + 3 is 0 so, z = 5 + 3

OpenStudy (anonymous):

8?

OpenStudy (goose.):

Yes :)

OpenStudy (goose.):

And as she was saying, you can check your answer after solving by plugging it into your variable.

OpenStudy (anonymous):

thanks

OpenStudy (goose.):

Glad to help! :)